4(x+3)-2(2)=1.5x+25.5
Answer by edjones(8007)

4(x+3)-2(2)=1.5x+25.5
4x+12-4=1.5x+25.5
4x-1.5x+8=1.5x-1.5x+25.2
2.5x+8-8=25.2-8
2.5x=17.2
2.5x/2.5=17.2/2.5
x=6.88
.
Ed